Solve the following linear programming graphically [5] Minimize: 𝑧 = 200π‘₯ + 500𝑦 Subject to the constraints: π‘₯ + 2𝑦 β‰₯ 10 3π‘₯ + 4𝑦 ≀ 24 π‘₯ β‰₯ 0; 𝑦 β‰₯Β 


1
Expert's answer
2021-10-08T09:40:48-0400

Minimize "\ud835\udc67 = 200\ud835\udc65 + 500\ud835\udc66"

subject to the constraints


"\begin{matrix}
 x+2y\geq10 \\
3𝑥 + 4𝑦 \leq 24 \\
x\geq0, y\geq0
\end{matrix}"

Find the point(s) of intersection


"y=-\dfrac{1}{2}x+5"

"y=-\dfrac{3}{4}x+6"

"x=0:"


"y=-\dfrac{3}{4}(0)+6, Point\ A(0,6)"

"y=-\dfrac{1}{2}(0)+5, Point\ C(0,5)"

"-\dfrac{1}{2}x+5=-\dfrac{3}{4}x+6"

"\dfrac{1}{4}x=1"

"x=4, y=3,Point\ B(4,3)"


Point "A(0,6):\ud835\udc67(0,6) = 200(0) + 500(6)=3000"


Point "B(4,3):\ud835\udc67(4,3) = 200(4) + 500(3)=2300"


Point "C(05):\ud835\udc67(0,5) = 200(0) + 500(5)=2500"


The function "z" has a minimum with value of "2300" at "(4,3)."